OpenStudy (anonymous):

Consider the line 2x-4y=8. What is the slope of a line perpendicular to this line? What is the slope of a line parallel to this line?

OpenStudy (alexandervonhumboldt2):

perpendicular lines have negative reciprocal slopes parallel line have same slopes

OpenStudy (anonymous):

Wait really, so what wuold I have to do?

OpenStudy (alexandervonhumboldt2):

first find slope of your line

OpenStudy (anonymous):

wait would it be for the perpindicular -- >-8 and for the paralell it would be 8?

OpenStudy (alexandervonhumboldt2):

2x-4y=8. -4y=8-2x y=(-2x+8)/(-4)=1/2x-2 the slope of your given line is 1/2

OpenStudy (alexandervonhumboldt2):

the slope if 1/2. the negative reciprocal to 1/2 is -2.

OpenStudy (anonymous):

so just to reiterate, the perpidicular would be -2 and the parellel would be the same?

TheSmartOne (thesmartone):

yes

OpenStudy (anonymous):

thanks other dude !

TheSmartOne (thesmartone):

no problem :P